OpenStudy (anonymous):

A delivery truck is transporting boxes of two sizes: large and small. The combined weight of a large box and a small box is 80 pounds. The truck is transporting 65 large boxes and 55 small boxes. If the truck is carrying a total of 4950 pounds in boxes, how much does each type of box weigh?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Lyrae

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well you need to solve simultaneously let l = weight of a large box and s = weight of a small box you know the weights l + s = 80 ( a large box and a small box weigh 80 lbs) 65l + 45s = 4950 combined weight of all the boxes so a simply way is substitution take the 1st equation and let s = 80 - l and substitute it into the 2nd equation 65l + 45(80 - l) = 4950 which becomes 65l - 45l + 3600 = 4950 20l + 3600 = 4950 so solve for l, then substitute it back into the very 1st equation to find s